OpenStudy (anonymous):

What is the equation, in standard form, of a vertical hyperbola with asymptotes at y + 8 = ±(x – 8)?

OpenStudy (amistre64):

you see that the slope of the asymps is 1 right?

OpenStudy (amistre64):

which means that a = 1 and b = 1 in the conventional hyperB equation

OpenStudy (amistre64):

\[\frac{x}{1}^2-\frac{y}{1}^2=1\] now the issue is the center; so where do these asymps cross at?

OpenStudy (amistre64):

y = +-x -+ 8- 8 y = +-x -16, +0 one has a yint of -16; the other at 0; so it looke to cross at y=-8

OpenStudy (anonymous):

Sorry, went to get coffee. I have( y-8)^2/81 - (x+8)^2/4 = 1

OpenStudy (amistre64):

where does 81 and 4 come from?

OpenStudy (amistre64):



OpenStudy (anonymous):

Sorry, I left off something form the original equation, ugghh. It is y+or - 2/9(x-8). So 81 comes from squaring the 9 and 4 comes from squaring the two

OpenStudy (amistre64):

ok .... then i will need to know what the whole question is. Since i am trying to answer the post itself :)

OpenStudy (amistre64):

asymps are: \(y=\pm\frac{2}{9}(x-8)\) right?

OpenStudy (anonymous):

What is the equation, in standard form, of a vertical hyperbola with asymptotes at y+8= plus/minus 2/9 (x-8)

OpenStudy (amistre64):

well, the point they share in common is: (8,-8) right?

OpenStudy (amistre64):

(x-8)^2 (y+8)^2 ------- () ------- = 1 a^2 b^2 if we open left to right along the x axis: (x-8)^2 (y+8)^2 ------- (-) ------- = 1 16 81 if we open up and down along the y axis: (y+8)^2 (x-8)^2 ------- (-) ------- = 1 16 81

OpenStudy (anonymous):

Ok. I had x as the leading term in the equation,but my choices only give y as a choice. Is that because it is a vertical hyperbole?

OpenStudy (anonymous):

Where did the 16 comes from? I though you would square the 2 and get 4

OpenStudy (amistre64):

.... yeah, i put 2 thoughts toghter; 2^2 = 4; but since its 4 ... i then requred it in my head

OpenStudy (amistre64):

resquared .. lol

OpenStudy (anonymous):

So you think the equation with 4 instead of 16 is correct?

OpenStudy (amistre64):

no, 2^2 = 4. 9^2 = 81 go with 4 and 81

OpenStudy (amistre64):

i did it right in the wolfram ....

OpenStudy (anonymous):

What is the worlfram, ha

OpenStudy (amistre64):

ummm nothing :) since i messed it up in there as well. go with the 4 and 81

OpenStudy (anonymous):

Is this a graphing program

OpenStudy (amistre64):

its a pretty neat graphing and other stuff engine yes

OpenStudy (anonymous):

Great! Is is free? I'm going to post another question if you want to help :)

OpenStudy (amistre64):

..... yes, its free